### v1

# Community-Cross-Instruct: Unsupervised Instruction Generation for Aligning Large Language Models to Online Communities

Abstract
Social scientists use surveys to probe the opinions and beliefs of populations, but these methods are slow, costly, and prone to biases. Recent advances in large language models (LLMs) enable creating computational representations or “digital twins” of populations that generate human-like responses mimicking the population’s language, styles, and attitudes. We introduce Community-Cross-Instruct, an unsupervised framework for aligning LLMs to online communities to elicit their beliefs. Given a corpus of a community’s online discussions, Community-Cross-Instruct automatically generates instruction-output pairs by an advanced LLM to (1) finetune an foundational LLM to faithfully represent that community, and (2) evaluate the alignment of the finetuned model to the community. We demonstrate the method’s utility in accurately representing political and fitness communities on Reddit. Unlike prior methods requiring human-authored instructions, Community-Cross-Instruct generates instructions in a fully unsupervised manner, enhancing scalability and generalization across domains. This work enables cost-effective and automated surveying of diverse online communities.
Community-Cross-Instruct: Unsupervised Instruction Generation for Aligning Large Language Models to Online Communities
Zihao He, Rebecca Dorn, Siyi Guo, Minh Duc Chu, Kristina Lerman USC Information Sciences Institute {zihaoh, rdorn, siyiguo, mhchu}@usc.edu, lerman@isi.edu
1 Introduction
Social scientists use surveys and focus groups to learn the opinions, needs and concerns of diverse populations. However, designing surveys and recruiting participants is a slow and costly process, limiting the utility of these instruments for probing public opinion. Surveys are prone to biases, such as the social desirability bias Gordon (1987), where respondents may alter their responses to sensitive questions to appear more socially acceptable Bergen and Labonté (2020), non-response bias Hill et al. (1997), where participants fail to answer questions, and self-selection bias due to the choices participants make to participate in the survey Heckman (1990). In addition, social stigmas may taint responses Goel and Salganik (2010), especially for hard-to-reach and marginalized groups.
Figure 1: Illustration of Community-Cross-Instruct to align an LLM to a community. (1) Open-ended instructions and multi-choice survey questions are generated by an advanced LLM from the community data. (2) A foundational LLM is aligned to community through instruction-tuning on the open-ended instructions. (3) The alignment of the finetuned LLM to the community is measured on the survey questions.
Recent breakthroughs in generative AI and especially large language models (LLMs) enable new capabilities for creating computational representations of human populations — their digital twins El Saddik (2018) — by ingesting vast textual data they create, for example, in online discussion forums. These LLM-based models generate human-like responses that mimic the language, communication styles, and attitudes of populations they are aligned to, allowing us to probe their worldviews, biases, and sentiments in a cost-effective and automated manner. Previous works have leveraged such LLM-based representations to mine opinions and learn political attitudes of online communities Jiang et al. (2020); He et al. (2024b). However, they finetune GPT-2 Radford et al. (2019) directly on textual data from communities, where the finetuned model is not able to answer questions that may be in different format, e.g., multiple choice, as it is not an instruction-tuned model.
To address these challenges, we introduce Community-Cross-Instruct: a fully unsupervised framework for aligning LLMs to online communities through instruction-tuning. The finetuned LLMs serve as digital twins of communities, which can be automatically surveyed to elicit their views (Figure 1).
Specifically, given a corpus of comments and submissions in different forums on Reddit, Community-Cross-Instruct uses an advanced LLM (Claude-3) to automatically curate (1) CommInst: a set of open-ended instructions (Fig. 2(a)) for community-specific instruction tuning, and (2) CommSurvey, a set of multi-choice survey questions (Fig. 2(b)). Each instruction in CommInst and each survey question in CommSurvey are paired with answers from different communities. We finetune foundational LLMs (GPT-3.5 or Llama-3) on CommInst, in order to align them to different communities, and evaluate the finetuned LLMs on CommSurvey to measure alignment. See Figure 3 for a high-level overview of the framework. Existing methods for self-supervised instruction generation rely on a seed set of instructions curated by domain experts Wang et al. (2022); Chen et al. (2024b), which is not generalizable. We overcome this bottleneck by incorporating the readily available community data into the instruction generation pipeline. Our Community-Cross-Instruct framework distills knowledge from communities using an advanced LLM to generate instructions for aligning foundational LLMs.
Our key contributions can be summarized as
  • We introduce Community-Cross-Instruct, a novel unsupervised framework for aligning foundational LLMs to online communities, by finetuning them on the automatically curated set of open-ended community-specific instruction-answer pairs (CommInst). The models’ alignment to communities is measured using another set of automatically generated multiple choice questions and answers (CommSurvey).
  • Using data from Reddit forums, we show that our method improves the fidelity of community representation (alignment) in two domains—politics and fitness—yielding significant alignment improvement over standard persona adaptation methods.
Our work highlights the potential of generative AI to help researchers gain insights from online communities. By leveraging LLMs to create digital twins of these communities, researchers can more accurately and efficiently understand the nuances of public opinion, attitudes, and behaviors. Our framework not only enhances the fidelity of community representation but also paves the way for more effective approaches to studying social phenomena in the digital age.
Clarification on contributions.
We aim to align LLMs to online communities through instruction-tuning, so that the aligned LLMs can be flexibly surveyed to probe the communities’ views on different topics. To prepare data for instruction tuning, we propose Community-Cross-Instruct, and the automatic creation of the open-ended instruction-answer pairs (CommInst) is the core contribution. However, in order to demonstrate that the finetuned LLMs are indeed aligned to communities and can be treated as their digital twins, we propose to automatically generate survey questions and answers (CommSurvey) to efficiently evaluate LLM alignment. This is a secondary contribution of this paper, and we do not claim that CommSurvey is the best way to measure alignment. Therefore, our main goal is to propose a method to prepare LLMs as a proxy for surveying different online communities, rather than automatically creating such surveys.
On the use of the term alignment.
Inspired by Santurkar et al. (2023), throughout this paper, we use alignment to refer only to the alignment of views and opinions of LLMs and humans.
2 Problem Definition
A topical domain (e.g., politics or fitness) includes $n$ communities {$C_{1}$, $C_{2}$, …., $C_{n}$} each with different views and beliefs. Members of each community $C_{i}$ collectively author text corpus $D_{i}$ (e.g., discussions on Reddit forums) expressing views and exhibiting behaviors. Our goal is to align an LLM $f$ to each community $C_{i}$ using its texts $D_{i}$, such that the aligned LLM $f^{\prime}_{i}$ learns the complex mindset of the community and responds to inputs in the community’s voice. By administering surveys to the aligned LLMs $\{f^{\prime}_{1},f^{\prime}_{2},...,f^{\prime}_{n}\}$, we obtain responses from different communities, thereby capturing their ideological differences.
To align an LLM $f$ to a community $C$, we finetune it on a set of demonstrations (instruction-answer pairs) Wang et al. (2023); Ouyang et al. (2022); Chen et al. (2024b) $I=\{(X_{j},Y_{j})\}_{j=1}^{m}$, where the instructions are open-ended questions probing the community’s views on different topics, and the corresponding answers are aligned with each community’s ideology. Figure 2(a) shows an example demonstration in the politics domain. We propose Community-Cross-Instruct (Fig. 3), a framework to automatically generate community-specific demonstrations $I$ with an advanced LLM $\hat{f}$ (Claude-3) based on the community’s text corpus $D$ and use these demonstrations to instruction-tune a foundational LLM $f$ (GPT-3.5 or Llama-3) through a process we call “Cross-Instruct”.
Figure 2: Example of (a) an instruction from CommInst-Pol and (b) a survey question from CommSurvey-Pol. The instruction/survey question is paired with answers from different communities.
Figure 3: Overview of Community-Cross-Instruct, with an illustrative example of the politics domain. (1) Data collection for each community within the desired domain. (2) BERTopic clusters the data and identifies prominent topics. (3) For each topic, the advanced LLM is prompted with (i) on-topic examples from each community and (ii) instruction-answer task definition and explanation. (4) Creation of (a) open-ended instruction-answer pairs and (b) multi-choice survey questions and answers, for fine-tuning and evaluating a sub-population representation foundational LLM respectively.
3 CommInst: A Collection of Community-Specific Instructions
To align LLMs to different communities within each domain, we curate a dataset CommInst for community-specific instruction tuning. CommInst-Pol and CommInst-Fit are the two subsets for the politics and fitness domains with 6,687 and 4,395 unique instructions respectively. Each instruction is an opinion-eliciting question about a topic relevant to the domain (Figure 2(a)), which appears in multiple demonstrations. Each demonstration’s instruction is paired with an open-ended answer from a different community.
3.1 Online Community Corpora Collection
Reddit is a vibrant social media platform hosting discussion forums (subreddits) on a wide range of topics Hofmann et al. (2022); Chen et al. (2024a). In this paper we focus on two domains: politics and fitness, which contain distinctive subreddits, or communities, with complex social dynamics. Political subreddits are valuable expressions of diverse public opinions and viewpoints. In the age of rising polarization, LLMs can help researchers track the complex evolving ideological landscape and effectively elicit public opinions. Meanwhile, the fitness subreddits feature a wealth of sensitive, health-related conversations, and the latest diet and fitness fads, often discussed using inscrutable insider jargon. We can track the direction of these discussions, identifying emerging trends, risks, and potential health misinformation. This insight allows public health officials to promptly address harmful health advice and intervene where necessary. Additionally, it provides an early warning system to detect the spread of dangerous health practices, ensuring that corrective measures can be implemented swiftly to protect community health.
We identify a set of representative subreddits for each domain based on personal knowledge and by querying ChatGPT. We manually aggregate and filter results, obtaining the following political online communities: r/Anarchism, r/Anarcho_Capitalism, r/Conservative, r/Libertarian, r/Republican, and r/democrats; For fitness, we investigate: r/Fitness, r/loseit, r/keto, r/xxfitness, and r/GettingShredded. More details of these communities are presented in Appendix A.1.
We collect comments and submissions from January 2019 to December 2022 and remove those that were deleted by the author or the moderator. The statistics of the collected comments and submissions are shown in Appendix A.1. To balance the number of comments and submissions across different communities, for each community, we keep a maximum of 140K comments and 30K submissions. Therefore, there are at most 170K texts, which makes up the text corpus $D$, including comments and submissions, for each community.
3.2 Topic Modeling
For a domain, denote the combined corpus from all $n$ communities by $\mathbb{D}=D_{1}\cup D_{2}\cup....\cup D_{n}$. We use BERTopic Grootendorst (2022) on $\mathbb{D}$ to identify topics $g$. More details about BERTopic can be found in Appendix B.1. After topic modeling, each text $d$ is assigned a topic $g(d)$. The number of topics detected in politics and fitness are 425 and 628 respectively. The top 10 most frequent topics in politics and fitness are shown in Appendix B.1 respectively. Politics topics include abortion, climate, guns, and racism. Fitness topics include body measures, keto diet, and weight lifting.
For each community $C_{i}$, we split the text corpus $D_{i}$ into smaller chunks $D_{i}=\{\hat{D}^{i}_{1},\hat{D}^{i}_{2},....\}$, where each chunk $\hat{D}$ consists of texts belonging to a single topic $g(\hat{D})$.
3.3 Open-ended Instruction-Answer Pair Generation
For a specific domain, starting with an empty demonstration pool, we randomly sample without replacement a text chunk $D_{i}$ for each topic $t_{k}\in T$ from each community. We include the sampled text chunks on topic $t_{k}$ from the $n$ communities into a prompt and then use the prompt to query the advanced LLM to generate open-ended instructions and answers on the topic. The prompting template is shown in Appendix B.2. We query claude-3-opus-20240229 using the Anthropic API. In the prompt, we first specify the topic by the topic keywords, and then the comments in the sampled chunks from different communities. Next we ask the model to generate instructions that can be answered based on the texts, and that the instructions should elicit different responses from the $n$ communities. Each generated instruction is paired with different answers from different communities (Figure 2(a)). It is then added to the pool as $n$ demonstrations. Note that some topics may be discussed more in certain communities than others, so some instructions may be paired with answers from less than $n$ communities. We iteratively repeat this process until there are no text chunks left for topic $t_{k}$ from any community, until we shift to the next topic $t_{k+1}$. After the iterative generation process, the instruction-answer pool will contain demonstrations covering various topics, which will use to finetune foundational LLMs. Please refer to Appendix A.2 for more details about CommInst.
For fair evaluating the finetuned LLMs, we split CommInst-Pol in two ways:
  • Random split. CommInst-Pol is randomly split into training and eval sets: CommInst-Pol-train and CommInst-Pol-eval, in a ratio of 85:15.
  • Topical split. We randomly select 10 topics in politics and collect demonstrations related to the topics into CommInst-Pol-eval-topic. The remaining demonstrations are collected as CommInst-Pol-train-topic. In this way, CommInst-Pol-train-topic and CommInst-Pol-eval-topic has no topic overlap.
We repeat the same process for CommInst-Fit.
4 CommSurvey: A Collection of Community-Specific Survey Questions
To measure the alignment of the finetuned foundational LLM to the community, we administer a survey of multi-choice questions to the model and evaluate the agreement between the model’s responses and the community members’ responses. However, manually designing surveys and collecting human responses from online communities is a non-trivial process, which is costly and time-consuming. Instead, we curate another dataset called CommSurvey for evaluating the alignment of the finetuned LLM’s views to the community using survey questions. CommSurvey consists of automatically generated multi-choice survey questions and answers, where each question is paired with answers from different communities (Figure 2(b)). CommSurvey-Pol and CommSurvey-Fit are the two subsets for politics and fitness respectively, consisting of 4,457 and 2,930 multi-choice questions respectively, covering a wide range of topics.
politics r/Anarchism
r/Anarcho_
Capitalism
r/Conservative r/Libertarian r/Republican r/democrats
0.75 0.60 0.75 0.45 0.75 0.65
fitness r/Fitness r/loseit r/keto r/xxfitness r/GettingShredded
0.75 0.70 0.60 0.50 0.65
Table 1: Agreement between human annotators’ survey responses and the advanced LLM (Claude-3) generated survey answers in different domains and subreddits, measured by accuracy.
4.1 Multi-choice Survey Generation.
We initialize an empty multi-choice survey pool. Then, following the open-ended instruction-answer pairs generation in §$\S$3.3, we iteratively generate multi-choice questions and answers (Figure 2(b)) using the same advanced LLM (Claude-3) and add them to the pool. We also explored other advanced LLMs, such as GPT-4 or Gemini, but none of them was able to generate high-quality questions or answers as Claude-3. The prompting template is shown in Figure 8. We assume that the answers generated by the advanced LLM are the “semi-ground truths”, and that they faithfully represent the views of the corresponding communities.
Note that there is a correspondence between instructions in CommInst and survey questions CommSurvey, as they are generated using nearly identical prompts. Following the correspondence, we split CommSurvey-Pol and CommSurvey-Fit following the same two ways for CommInst. Due to the correspondence, there is no information leakage between CommInst-Pol-train and CommSurvey-Pol-eval, or between CommInst-Pol-train-topic and CommSurvey-Pol-eval-topic.
4.2 Validating Generated Survey Answers
We verify the faithfulness of the “semi-ground truths” for the communities via human annotation. For each domain, we randomly sample 20 questions from the survey (see Fig. 9 and Fig. 10 in the Appendix) and ask two expert annotators to answer them from the perspective of the subreddit. We also tried directly posting the surveys to the subreddits and having community members answer them. However, our posts were immediately removed by moderators and our accounts were banned from certain subreddits. This further indicates the difficulty of directly surveying online communities.
[…]
5 Experiments
We demonstrate the effectiveness of the pipeline on the Reddit politics and fitness domains. We align LLMs to Reddit communities by finetuning them on relevant demonstrations from CommInst. After finetuning, we administer surveys from CommSurvey and evaluate the finetuned LLMs by comparing their responses to the “semi-ground truths”, in order to measure alignment.
5.1 Experimental Setup
We experiment with two foundational LLMs – GPT-3.5 and Llama-3.
Finetuning.
We finetune the foundational LLMs on CommInst-Pol-train and CommInst-Fit-train. We use the OpenAI finetuning API to finetune gpt-3.5-turbo-0125 for 1 epoch, as we find that the loss converges within 1 epoch. The batch size is automatically determined by the API, and the finetuning takes ~40mins. For Llama-3, the implementation is based on LlamaFactory Zheng et al. (2024), and we finetune Llama-3-8B-Instruct with batch size 16 for 2 epochs on 4 H100 GPUs. The finetuning takes ~10mins.
Measuring Alignment.
We administer survey questions in CommSurvey-Pol-veal and CommSurvey-Fit-veal to the finetuned foundational LLMs. In the prompt, we provide the question and the four options verbatim, and then add the instruction “Select only one answer by stating either A, B, C, or D. Do not provide any additional explanation or rationale for your choice.”, for the sake of easier matching of the model’s response to the option. For GPT-3.5, the temperature is set to 0.9 and the model provides 5 generations for each question. For Llama-3, temperature is set to 0.9, $top\_p$ is set to 0.9, and the model provides 10 generations for each question. The majority vote from the different generations for each question is taken. We compare the model responses to the “semi-ground truths” and report the accuracy as a measure of alignment.
5.2 Baselines
Vanilla LLMs.
We administer survey questions to the vanilla unfinetuned GPT-3.5 and Llama-3 and evaluate their answers.
Vanilla LLMs+Steering.
When prompting the vanilla language models to answer survey questions, we steer them to mimic the community, by specifying in the prompt that “Select only one answer that best aligns with the opinions of members from subreddit r/[subreddit]”. Steering the vanilla LLMs can nudge them to respond for the community. However, it is important to note that steering only applies to predefined communities that are developed via a predefined tag, such as subreddits, where we can easily specify community names in the prompt to steer the LLMs. Although in this paper we focus on communities associated with online forums, Community-Cross-Instruct generalizes to other online communities, including organically formed communities in the retweet network Chu et al. (2024) or the news co-sharing network Mosleh et al. (2021); He et al. (2024b), as long as their texts are readily available. It is non-trivial to summarize the organically-formed communities concisely in language, but we can still align LLMs to them via Community-Cross-Instruct.
5.3 Main Results
We compare survey answers generated by the finetuned LLMs to the “semi-ground truth” and report the accuracy in Figure 4 as a measure of alignment with the corresponding community. In most cases, steering LLMs towards specific communities creates better alignment by relying on the models’ pre-existing knowledge about the communities. However, such pre-knowledge may be incomplete, inaccurate, or outdated. Therefore, by finetuning LLMs on instructions distilled from the text corpora of communities, Community-Cross-Instruct outperforms steering and achieves the best results, demonstrating its effectiveness in aligning LLMs to different communities.
(a) politics
(b) fitness.
Figure 4: Alignment of finetuned LLMs to the corresponding communities measured by accuracy, by comparing the LLMs’ survey responses to the “semi-ground truths”.
The base models show worse alignment with r/Conservative and r/Republican than other Reddit communities. This observation conforms to findings by previous works Santurkar et al. (2023); He et al. (2024a); Chen et al. (2024b) that LLMs tend to display a liberal bias and poor alignment with conservatives. After explicitly steering them towards specific communities, their alignment with the two subreddits significantly increases, but it is still lower than that with other communities. Finetuning the LLMs with Community-Cross-Instruct further improves the alignment, although it is still lower than for other communities. This observation highlights the risks of deploying LLMs in political scenarios due to the entrenched political biases.
For the r/Anarchism subreddit, GPT-3.5-CrossInst achieves similar alignment as the steering baseline. We suspect that this is because views about anarchism are captured by the pre-existing knowledge of GPT-3.5. Therefore, infusing such knowledge again by Community-Cross-Instruct barely leads to any gain.
For the r/Fitness and r/xxfitness communities, steering both LLMs worsens the alignment compared to the vanilla model. r/Fitness is a general fitness subreddit and r/xxfitness is a fitness community for women, both of which may have a broader and more diverse range of opinions compared to other subreddits. Steering the models to align with a singular “community opinion” might oversimplify the nuanced perspectives within these fitness communities, leading to decreased alignment accuracy. However, Community-Cross-Instruct effectively distills knowledge from the community corpora during finetuning. On r/GettingShredded, Community-Cross-Instruct is outperformed by the steering baseline for GPT-3.5, but not Llama-3. r/GettingShredded has the least amount of comments and posts, producing much fewer demonstrations for finetuning (Appendix 3). The insufficient data might be less effective for finetuning a larger LLM like GPT-3.5111We assume that GPT-3.5 has more than 8B parameters., leading to inferior performance.
5.4 Out-of-Topic Generalizability
Our final goal is to create LLMs aligned to different communities, which can be used to answer any survey question from the perspective of the community. In the real world, the surveys may contain questions covering topics that do not appear in the finetuning data CommInst. In this study, we finetune Llama-3 on CommInst-Pol-train-topic, and evaluate it on CommSurvey-Pol-eval-topic, to make sure that there is no overlap in topic coverage. The results in Fig. 5 show that Community-Cross-Inst consistently outperforms both baselines, demonstrating its superiority in aligning with views of different communities, and generalize the learned views to unseen topics.
Figure 5: Cross-topic alignment of finetuned LLMs to the corresponding communities. The LLMs are finetuned on demonstrations and evaluated on survey questions on different topics.
6 Related Works
Self-Improved Instruction Tuning. A survey of instruction-tuning approaches Zhang et al. (2023) outlines several methods for models to autonomously self-improve their instruction set. These include generating instructions for pre-existing texts Li et al. (2023), eliciting interaction between different model iterations Chen et al. (2024c), and bootstrapping from an existing instruction set to generate new ones Wang et al. (2022). Additionally, a human-in-the-loop framework builds upon self-generalization by iterating between human and machine-generated instructions Guo et al. (2024). Despite these advances, these strategies require manual input. In this work, we build on self-instruct, a method to enhance instruction tuning by eliciting synthetic model instruction generations. We alter the original framework by removing the required set of manually written seed instructions Wang et al. (2022). This work significantly contributes to the field by offering a scalable and fully autonomous solution to instruction tuning, paving the way for more adaptive and intelligent models.
Community Alignment. Aligning with the dialects and behaviors of online communities is crucial for accurately modeling subpopulations Dorn et al. (2023, 2024). Subpopulation representative models (SRMs) Simmons and Hare (2023) can be used to emulate some characteristics of a particular subpopulation, particularly as LLMs can provide fine-grained, demographically-correlated outputs Argyle et al. (2023a). Using self-instruct with questions from the American Trans Panel as the seed instruction set, LLMs can be ideologically manipulated to reflect particular political opinions Chen et al. (2024b). To learn about partisan communities, Jiang et al. (2022) repeatedly prompt GPT 2 with tweets authored by prominent community members to generate synthetic responses to election survey questions. Argyle et al. (2023b) find that exposing GPT 3 to thousands of socio-demographic backstories leads the model to obtain a complex understandings of sociological concepts. ‘Media diet models’ train BERT using MLM on news articles, evaluating the alignment of model predictions with a survey of community members who follow the media Chu et al. (2023). Feldman et al. (2022) expose GPT 3 to Yelp reviews to expand insights past sentiment information. Prompting language models with (survey question, individual ID, survey year) triplets from the Americans in the General Social Surveys has helped social scientists impute missing survey data and predict opinions Kim and Lee (2023). These techniques rely on manually created datasets at least as a seed dataset to align the models with the desired communities, limiting SRBMs to high-resource environments.
Surveys. Social scientists use surveys to systematically collect data from populations to characterize their beliefs, attitudes, opinions, and behaviors. While designing the survey, researchers consider the order, wording, and format of questions (open-ended vs multiple-choice) to increase clarity and ensure the validity and reliability of the survey data. Survey design also aims to minimize biases, such as the non-response bias Hill et al. (1997), fatigue due to overly long surveys Choi and Pak (2005), social desirability bias Gordon (1987) when asking survey participants questions about sensitive topics (due to the desire to conform to perceived social expectations), or those involving social stigmas Goel and Salganik (2010). These issues, as well as the variety of existing biases, call for new approaches to survey design.
7 Conclusion
In this paper, we introduced and explored Community-Cross-Instruct, a fully automated framework to represent online communities and elicit their views. By automating the process of surveying diverse online communities, Community-Cross-Instruct offers a cost-effective and efficient alternative to traditional survey methods used by social scientists. Through finetuning LLMs based on community-specific instructions and answers, Community-Cross-Instruct enables the generation of accurate responses that align with the beliefs and perspectives of different online communities. This innovative framework has demonstrated its effectiveness in representing various communities, including political and fitness forums on Reddit. Community-Cross-Instruct opens up new possibilities for researchers to gain insights into the diverse views and opinions present in online communities, paving the way for deeper understanding and analysis of digital societies. Moving forward, we will apply the framework to model organic online communities, such as those in the retweet network.
Limitations
Identical LLM for Generating both Datasets.
We use Claude-3 for generating both CommInst for finetuning LLMs and CommSurvey for evaluating LLM alignment, as we couldn’t find a different advanced LLM that was able to generate high-quality data as Claude-3. This reliance on a single LLM introduces a limitation, as it may lead to a bias where the evaluation dataset (CommSurvey) is inherently aligned with the model used for instruction generation (CommInst), potentially overestimating the alignment accuracy of the finetuned models. To address this limitation, we plan to incorporate a diverse set of advanced LLMs as they become available in future iterations of our framework. Additionally, we will include human-in-the-loop validation to ensure the generated datasets maintain high quality and representativeness, mitigating any biases introduced by the single LLM dependency. Furthermore, we will explore cross-validation techniques and third-party evaluations to benchmark the performance of our finetuned models, ensuring robustness and generalizability of our results beyond the influence of a single LLM.
Group Approximation.
To approximate group-level behavior, community members represented in the minority are inherently excluded. Further, measuring group alignment using a single-answer multiple choice questionnaire does not account for bimodal distribution. We hope to build on this work by experimenting with survey designs that account for more diversity in communities.
Hallucination Potential.
Adapting models to communities poses a risk of language models hallucinating or providing misinformation in their community representation. We hope to build upon this work with in-depth experiments on model hallucination and misrepresentation in the subgroup representation task.
Prompt Perturbations.
LLM responses are sensitive to slight changes in prompts Salinas and Morstatter (2024). In this work, we work primarily with one prompt for instruction generation. We would be interested to see how model’s generated instructions shift with different prompting schemas.
Ethics Statement
Users might not be informed of how their data (reflected in their posts) are used to produce digital twins that mirror their voices and the purpose of the survey constructed based on their data. Furthermore, automatically surveying communities can reveal unconsented insights of certain individuals or groups of people online. To address these ethical concerns, we will implement several measures. Firstly, we will ensure that all data used for creating digital twins is anonymized, stripping any personally identifiable information to protect user privacy. Secondly, we will seek to aggregate data in a manner that focuses on community-level insights rather than individual-level analysis, thereby reducing the risk of unconsented personal exposure. Thirdly, we will obtain explicit consent from users where possible, clearly communicating the purpose and scope of our research. Lastly, we will adhere to ethical guidelines and institutional review board (IRB) requirements to ensure that our methods respect the privacy and rights of the individuals and communities involved.
Appendix A Data Details
A.1 Subreddits
We briefly introduce each subreddit in Table 2. Statistics of data from different subreddits are shown in Table 3.
Subreddit Description
r/Anarchism
r/Anarchism is a forum with 268K members for discussing anarchism, assuming
participants are anarchists, allies, or learners.
r/Anarcho_Capitalism
r/Anarcho_Capitalism is a subreddit with 197K members dedicated to discussing
free-market capitalist anarchism and related topics, advocating for a society where
voluntary interactions enhance liberty and opportunity for all.
r/Conservative
r/Conservative, with 1.1M members, offers the largest space on Reddit for fiscal
and social conservatives to explore and discuss political and cultural issues from
a distinctly conservative perspective.
r/Libertarian
r/Libertarian is a forum with 506K members focused on libertarianism and the
principles of a free society, explicitly rejecting anti-libertarian ideologies
like communism.
r/Republican
r/Republican is a partisan subreddit with 191K members specifically designed
as a space for Republicans to discuss issues among themselves.
r/democrats
r/democrats is a subreddit with 438K members dedicated to supporting the
Democratic Party, offering daily updates, policy analysis, and opportunities to
help elect Democrats at all levels.
r/Fitness
r/Fitness is the largest community dedicated to achieving physical fitness
goals with 12M members
r/loseit
r/loseit is a supportive community with 4M members for individuals of all
sizes to discuss healthy and sustainable weight loss methods
r/keto
r/keto is a community of 3.6M members for sharing experiences and advice
about the low-carb Ketogenic diet, which supports a range of health
conditions from diabetes to epilepsy
r/xxfitness
r/xxfitness is a community of 2.2M members for female and gender
non-binary/gender non-conforming Redditors to discuss fitness
r/GettingShredded
r/GettingShredded is a community of 162K members for anyone aiming to lose
body fat, improve their physique, and boost confidence through healthy fat loss
and strength training discussions.
Table 2: Reddit forums used in this study. Descriptions of (top) politics and (bottom) fitness subreddits.
A.2 Statistics of CommInst
Table 3 describes the number of demonstrations in each subreddit. The variance is due to their different topic coverage. Figure 6 presents the distribution of instruction lengths and answer lengths in CommInst-Pol and CommInst-Fit.
Politics
r/Anarch. r/Anarcho. r/Conserv. r/Democr. r/Libert. r/Republic.
Comments 182,192 589,354 2,391,745 142,274 1,950,089 140,572
Submissions 31,567 41,264 356,842 90,001 40,967 48,808
Demonstrations 2,430 3,027 2,943 3,159 2,928 2,574
Fitness
r/Fitness r/loseit r/keto r/xxfitness r/GettingShredded
Comments 907,839 967,645 603,920 420,082 78,053
Submissions 2,430 3,027 2,943 3,159 2,928
Demonstrations 3,438 2,793 1,716 2,883 1,215
Table 3: Dataset size measured by number of comments and submissions for each community in political and fitness communities, and the number of learned open-ended instruction-answer pairs (demonstrations).
(a) Instructions in CommInst-Pol
(b) Instructions in CommInst-Fit
(c) Answers in CommInst-Pol
(d) Instructions in CommInst-Fit
Figure 6: Distribution of instruction and answer lengths in CommInst-Pol and CommInst-Fit.
Appendix B Implementation Details
B.1 Topic Modeling with BERTopic
Following the pipeline of BERTopic Grootendorst (2022), we first obtain the text embeddings using all-mpnet-base-v2 Song et al. (2020). The embedding dimensions are reduced with UMAP, with n_neighbors 15 and n_components 5. Then the embeddings are clustered using HDBSCAN, with min_cluster_size 50. After fitting the model, we remove texts that are not assigned any topics. Table 4 and Table 5 present the top-10 most frequent topics in politics and fitness.
idx Topic Keywords Frequency
1 covid, vaccinated, masks, vaccines, healthcare, deaths, flu, pandemic, coronavirus, cdc 48,787
2 abortion, fetus, abortions, roe, body, choice, unborn, rights, parenthood, cells 11,256
3 climate, change, solar, emissions, pipeline, fuels, coal, power, supply, grid 9,035
4 guns, firearms, shootings, firearm, amendment, laws, armed, rifles, militia, nra 8,833
5 police, officer, officers, gun, force, law, crime, video, training, situation 7,934
6 kamala, woman, pelosi, vp, hillary, president, herself, biden, party, with 7,879
7 racist, racism, whites, racists, racial, minorities, democrats, privilege, whiteness, discrimination 7,785
8 libertarian, libertarians, libertarianism, party, government, left, ideology, vote, political, conservative 6,666
9 anarchism, anarchy, anarchist, anarchists, anarcho, hierarchies, communism, ideology, government, system 6,311
10 biden, bernie, obama, candidate, voters, debate, election, democrats, hillary, sanders 5,768
Table 4: Top-10 most frequent topics in politics. Each topic is represented by the topic-10 keywords.
idx Topic Keywords Frequency
1 lbs, bmi, down, cw, gained, progress, healthy, tips, weeks, started 18,296
2 keto, ketosis, carbs, diet, 20g, protein, macros, started, calories, ketogenic 13,078
3 leggings, clothes, wear, pants, shorts, bras, shirts, clothing, dress, lululemon 7,063
4 motivation, journey, progress, yourself, change, motivated, success, start, proud, back 5,142
5 water, scale, weighing, retention, trend, normal, plateau, fluctuates, gain, weekly 4,880
6 sleep, morning, hours, work, gym, motivation, schedule, waking, exercise, routine 4,486
7 she, people, overweight, weight, obese, comment, healthy, skinny, talk, shaming 4,192
8 bf, percentage, fat, bodyfat, body, lean, bulk, estimates, lbs, pic 3,930
9 reps, sets, program, max, amrap, lifts, ranges, bench, 5x10, hypertrophy 3,788
10 shoulder, elbow, injury, arm, knees, physio, exercises, bicep, curls, heal 3,775
Table 5: Top-10 most frequent topics in fitness. Each topic is represented by the topic-10 keywords.
B.2 Prompting Template for CommInst
The prompting template for generating the open-ended instruction-answer pairs in CommInst is shown in Figure 7.
Figure 7: Template prompt to generate CommInst.
B.3 Prompting Template for Generating CommSurvey
The prompting template for generating the multiple choice instruction-answer pairs in CommSurvey is shown in Figure 8.
Figure 8: Template prompt to generate CommSurvey.
Appendix C Survey Sample
Samples of the surveys provided to human annotators for generating “semi-ground truth data” are provided in Figures 9 and 10 (political and fitness domains, respectively).
Figure 9: Sample of the multiple choice survey generated for human evaluation in the politics domain.
Figure 10: Sample of the multiple choice survey generated for human evaluation in the fitness domain.